Please help, I am stuck. Thank you
The length of the base of an isosceles triangle is one fourth the length of one of its legs. If the perimeter of the triangle is 90 cm, what is the length of the base? I would really appreciate any help you could give.

Answer by checkley71(8403)

ISOSCELESE TRIANGLES HAVE 2 EQUAL SIDES.
BASE=X/4 WHERE X=LENGTH OF THE EQUAL SIDES.
2*X+X/4=90
2X+X/4=90
(4*2X+X)/4=90
(8X+X)/4=90
9X=360
X=360/9
X=40 ANSWER FOR THE LENGTH OF THE EQUAL SIDES.
BASE=40/4=10 LENGTH OF THE BASE.
PROOF
2*40+10=90
80+10=90
90=90